# Quantized distributed Nash equilibrium seeking under DoS attacks

### v2

@@ I Introduction
Game theory has attracted tremendous attention owing to its advances in decision making. It has great potentials in the applications such as wireless communication networks, defending cyber attacks and cloud computing [1, 2]. Nash equilibrium (NE) seeking is one of the most important problems in non-cooperative games. Recently, distributed NE seeking with partial-decision information becomes a new thriving direction, in which players do not need to have direct access to the information of all the other players [3, 4, 5].
@@
Under an unreliable network with cyber attacks, players’ actions may not reach a NE [18, 19]. Our paper considers one of the most common cyber attacks: Denial-of-Service (DoS) attacks. The intention of DoS attackers is to maliciously jam the networks. Communication failures induced by DoS can exhibit a temporal profile quite different from the one induced by genuine packet losses. Particularly, packet dropouts induced by DoS need not follow a given class of probability distributions [20]. Therefore, the techniques relying on probabilistic arguments for random packet losses may not be applicable in the context of DoS attacks. In the past decade, control over DoS-corrupted networks has been widely studied [21, 22]. The paper [22] proposed a new distributed resilient control method to address the output regulation problem under DoS attacks, without pre-knowing the dynamics of the leader. Since DoS attacks can induce communication failures, the existing distributed NE seeking strategies and results, e.g., in [11, 6, 12, 23], may not always hold.
@@
When quantization and DoS attacks coexist, NE seeking problems are challenging because one needs to ensure that quantizers should not saturate under DoS attacks. For example, under DoS attacks, states could diverge. If a quantization scheme has only zooming-in capabilities as in [12], quantizer saturation will occur under DoS attacks. To mitigate quantizer saturation problems in the presence of random or malicious packet losses, one can develop quantization strategies with both zooming-in and zooming-out/holding capabilities, for centralized systems [24, 25] and multi-agent systems [26, 27]. For instance, the works [24, 25] developed zooming-in and out quantization strategies to stabilize centralized systems under random and DoS-induced packet losses, respectively. In [26, 27], the authors proposed zooming-in and out, and zooming-in and holding strategies, respectively, to tackle quantized consensus problems under DoS attacks. The approach in our paper is inspired by the ideas of consensus based NE seeking in [6], dynamic quantized consensus under DoS attacks in [26, 27] and quantized consensus in the seminal papers [28, 29].
@@
Our paper deals with quantized distributed NE seeking problems under DoS attacks. We are interested in three questions: 1. how to design a new quantization scheme realizing NE seeking and preventing quantizer saturation under DoS attacks; 2. how robust is the quantized distributed NE seeking algorithm; 3. under DoS attacks, what is the lower bound of the number of quantization levels preventing quantizer saturation. Different from the Lyapunov-based approaches [11, 12], we develop a new path to answer the above questions, which is aligned with the approaches for studying quantized consensus [28, 26, 29, 27]. We mention that such a new development is challenging from the technical viewpoint because NE seeking problems involve nonlinear functions, under which some key techniques for linear-system consensus in [28, 26, 29, 27] are not applicable. The main contributions of this paper are listed as follows:
@@ II-B Time-constrained DoS
For more discussions regarding Assumptions 3 and 4, we refer the readers to [33] and [30]. Note that Assumptions 3 and 4 were proposed in [33] investigating centralized systems. Later, they have been widely adopted in the works of multi-agent systems, e.g., [34, 35, 36].
@@
[25] Consider the DoS attacks characterized by Assumptions 1 and 2 and the network sampling period $\Delta$. If $11/T+\Delta/\tau_{D}<$ 1, then $T_{S}(0,k\Delta))\geq\left(1-1/T-\Delta/\tau_{D}\right)k-\frac{\kappa+\eta\Delta}{\Delta}.$  $\blacksquare$
@@ II-C Quantizer
The limitation of bandwidth implies that transmitted signals are subject to quantization. Let $b\in\mathbb{R}$ be the original scalar signal before quantization and $q_{R}(\cdot)$ be the quantization function for scalar input values as
@@ II-D Control objectives
    Design a quantized distributed NE seeking controller and find a $R$ (or equivalently $(2R+1)/2$) such that the quantizer (8) is not saturated for all $k$, i.e., $|b|\leq(2R+1)/2$.
@@ III-A Control architecture
with $\theta(0)=\theta_{0}\in\mathbb{R}_{>0}$. In this paper, we select the zooming-in factor $\gamma_{1}<1$ and the “zooming-out factor” $\gamma_{2}=1$. Since $\gamma_{2}=1$, “zooming-out” also implies “holding”. The reasons of such $\gamma_{1}<1$ for zooming-in and $\gamma_{2}=1$ for holding will be specified later in Remark 3. Please note that acknowledgments are not necessary because each player is able to passively know the status of DoS at $k$ by whether receiving the neighbors’ transmissions. Thus, $\theta(k)$ can be updated locally. Since each player has the identical initial condition $\theta_{0}$, they can locally check $k\in H_{q}$ or $k\notin H_{q}$, one can verify that $\theta(k)$ is synchronized in all the players.
In (17), the update of $\theta(k)$ has two modes. The first mode is for NE seeking and the second is particularly for mitigating the influence of DoS attacks. Under DoS attacks, $y_{ij}(k+1)-\hat{y}_{ij}(k)$ and $x_{j}(k+1)-\hat{x}_{j}(k)$ in (14) can diverge [27] or be invariant [26]. If $\theta(k)$ always zooms in by only utilizing $\gamma_{1}$ as in [12], then as $\theta(k)$ approaches zero, $|(y_{ij}(k+1)-\hat{y}_{ij}(k))/\theta(k)|$ and $|(x_{j}(k+1)-\hat{x}_{j}(k))/\theta(k)|$ in $Q_{R}(\cdot)$ can approach infinity. Eventually, quantizer saturation will occur, e.g., $|(y_{ij}(k+1)-\hat{y}_{ij}(k))/\theta(k)|>(2R+1)/2$. Quantizer saturation will also occur in general if the players transmit quantized $y_{ij}(k+1)/\theta(k)$ to other players instead of following the mechanism in (13), even without DoS. This is because the growing $y_{ij}(k+1)/\theta(k)$ (due to $\theta(k)\to 0$) will exceed the range of any finite quantizer $q_{R}(\cdot)$. The same problem also exists for (14).
@@
By (11)–(17), one can see that the NE seeking controller has two modes depending on the status of DoS attacks. We propose such a switched-type controller for mitigating the influence of DoS attacks. The laws for quantized NE seeking in (12)–(17) are partially inspired by the approaches in quantized consensus [29, 28, 27, 26]. So will be the subsequent analysis. However, as will be shown later, due to the presence of nonlinear function in (11), the detailed technical analysis is different from quantized consensus.
@@ III-B Quantized system dynamics
Case 1) To obtain (22), it is important to note the following dynamics for Case 1): $u_{i}(k)=-\delta\frac{\partial f_{i}}{\partial x_{i}}(x_{i}(k),y_{-i}(k))$ by (11) and hence $x_{i}(k+1)=x_{i}(k)-\delta\frac{\partial f_{i}}{\partial x_{i}}(x_{i}(k),y_{-i}(k))$, $\sigma(k)=h$ by (12b) and hence $y_{ij}(k+1)=y_{ij}(k)-\sigma(k)(\sum_{l=1}^{N}a_{il}(\hat{y}_{ij}(k)-\hat{y}_{lj}(k))+a_{ij}(\hat{y}_{ij}(k)-\hat{x}_{j}(k)))$ by (12a), $\hat{y}_{ij}(k+1)=\hat{y}_{ij}(k)+\theta(k)\hat{Q}_{ij}(k+1)$ by (13a), $\hat{x}_{j}(k+1)=\hat{x}_{j}(k)+\theta(k)\hat{Q}_{j}(k+1)$ by (14a). Then, one can obtain the iterative equations of $\bar{y}(k)$, $x(k)$, $e^{x}(k)$ and $e^{y}(k)$ by the dynamic above and the definitions in (18)–(21).
@@
Case 2) To obtain (24), one should note that: $u_{i}(k)$, $x_{i}(k+1)$, $y_{ij}(k+1)$ follow the same dynamics as in Case 1) because of $k\notin H_{q}$. However, one has $\hat{y}_{ij}(k+1)=\hat{y}_{ij}(k)$ by (13a) and $\hat{x}_{j}(k+1)=\hat{x}_{j}(k)$ by (14a) because of $k+1\in H_{q}$.
@@
$e^{y}(k+1)=Ge^{y}(k)-hA_{0}(\mathbf{1}_{N}\otimes e^{x}(k))-hS\bar{y}(k)$ (24d)
Case 3) To obtain (25), one has $u_{i}(k)=0$ by (11) and hence $x_{i}(k+1)=x_{i}(k)$, $\sigma(k)=0$ and hence $y_{ij}(k+1)=y_{ij}(k)$ by (12) because of $k\in H_{q}$; $\hat{y}_{ij}(k+1)$ and $\hat{x}_{j}(k+1)$ follow the same dynamics as in Case 2) because of $k+1\in H_{q}$.
@@
$e^{y}(k+1)=e^{y}(k)$ (25d)
Case 4) To obtain (26), first note $x_{i}(k+1)=x_{i}(k)$ and $y_{ij}(k+1)=y_{ij}(k)$ due to $k\in H_{q}$ (see Case 3). Moreover, one should also note $\hat{y}_{ij}(k+1)=\hat{y}_{ij}(k)+\theta(k)\hat{Q}_{ij}(k+1)$ by (13a), $\hat{x}_{j}(k+1)=\hat{x}_{j}(k)+\theta(k)\hat{Q}_{j}(k+1)$ by (14a) due to $k+1\notin H_{q}$.
@@
$e^{y}(k+1)=e^{y}(k)-\theta(k)Q_{R}\left(\frac{e^{y}(k)}{\theta(k)}\right)$ (26d)
It is clear that if the network is free from DoS attacks, Case 1) is sufficient for analyzing the quantized distributed NE seeking problem. The presence of DoS attacks complicates the NE seeking problem by introducing Cases 2)-4). Consequently, the analysis becomes more challenged due to Cases 2)-4) as shown in the proofs of Lemma 3 and Theorem 1 later. Compared with quantized consensus of linear systems, the challenge is raised by the involvement of the nonlinear function $P(\eta(k))$.
@@ IV-A Analysis of zooming-in and holding quantization
We follow the approach aligned with that for quantized consensus under data rate constraints [29, 28, 26, 27] and define
@@
Note that a similar assumption on the initial condition is also required in the literature of quantized consensus [28, 29]. For quantized NE seeking problems, one also needs the assumption $C_{x^{*}}\geq|x^{*}_{i}|$. They both are required for calculating an upper bound of actions and preventing quantizer saturation at the initial time. We refer the readers to Remark 1 in [12] for more detailed reasons.
@@
in which $l:=\max\{l_{i}\}_{i\in\mathcal{N}}$. If $\|\xi^{y}(s_{g})\|_{\infty}\leq 1/2\gamma_{1}$ and $\|\xi^{x}(s_{g})\|_{\infty}\leq 1/2\gamma_{1}$ for $g=-1,0,\cdots,r-1$, then
@@
2) The purposes of Lemma 3 are two-fold: Specify the zooming-in parameter $\gamma_{1}$ and provide an upper-bound on $\|\pi(s_{r})\|$. Note that $\gamma_{1}\in\mathbb{R}_{<1}$ is necessary for the success of NE seeking even without DoS attacks. The result $\|\pi(s_{r})\|<C$ is also necessary since we will use it to show that $\|\pi(\cdot)\|$ is upper bounded for all $k$. Then one can obtain the main results: Quantizer unsaturation at all $k$ and NE seeking as $k\to\infty$.
@@ IV-B Quantizer unsaturation and convergence to the NE
Proof. First, we present the basic ideas of the proof. For result i), its proof is conducted by induction: If the quantizer is not saturated at $s_{p}$ ($p=-1,0,\cdots,r$ if DoS is present at $k=0$; $p=0,\cdots,r$ if DoS is absent at $k=0$), then it will not saturate for the transmissions at $k\in(s_{r},s_{r+1}]$. This also manifests that the quantizer is not saturated at $s_{r+1}$, which implies that the quantizer will not saturate for $k\in(s_{r+1},s_{r+2}]$. By induction, one can verify that the quantizer is not saturated at any $k$ in $s_{l}\bigcup_{r\in\mathbb{Z}_{\geq l}}(s_{r},s_{r+1}]$ ($l=-1$ if DoS is present at $k=0$; $l=0$ if DoS is absent at $k=0$), which are essentially all the transmission attempts. For obtaining result ii), it is sufficient to show that $\|\pi(k)\|$ is upper bounded and the scaling function $\theta(k)\to 0$ by the fact $\left\|\!\begin{bmatrix}\|\bar{y}(k)\|\\ \|x(k)-x^{*}\|\end{bmatrix}\!\right\|=\theta(k)\|\pi(k)\|$. In particular, the scaling function $\theta(k)\to 0$ if DoS attacks satisfy $11/T+\Delta/\tau_{D}<$ 1.
@@
where the first inequalities in (IV-B) and (IV-B) are obtained by Case 4.
@@
Step 5. In the following, we prove result ii). By (27) and (-B) (in Appendix-A), one can verify that $\|\pi(s_{r}+1)\|$ is upper bounded. Then by the first two equations in Cases 3) and 4), one can obtain $\|\pi(s_{r}+m)\|=\|\pi(s_{r}+1)\|$ for $m=2,3,\cdots$. Thus, one can see that $\|\pi(\cdot)\|$ is upper bounded at $s_{r}$ and $s_{r}+m<s_{r+1}$ ($m=1,2,\cdots$), which implies that $\|\pi(\cdot)\|$ is upper bounded at all $k$. By the transformation in (27) and the scaling function in (17), one can obtain
@@
By Theorem 1, it is clear that if the amount of DoS attacks characterized by $1/T+\Delta/\tau_{D}$ is smaller than 1, the actions of players converge to the NE under quantized signals. As indicated in [30], this is the best robustness one can achieve in the sense that if $1/T+\Delta/\tau_{D}\geq 1$, DoS attacks can be constantly present or fully synchronized with all the transmission attempts. When the communication links among the players are totally disabled, then it is impossible to solve the distributed NE seeking problem for any controller, to the best of our knowledge.
@@
Our result regarding zooming-out in dynamic quantization is consistent with those of quantized consensus [27] and quantized stabilization of centralized systems [24, 25]. In [27], the agents have a general form (e.g., $x_{i}(k+1)=Ax_{i}(k)+Bu_{i}(k)$) and hence the zooming-out factor $\gamma_{2}$ in [27] depends on the unstable eigenvalues of $A$. In our paper, the players take the form of an integrator or equivalently $A$, as a special case, is an identity matrix. Then the zooming out factor $\gamma_{2}$ takes the value 1. It is also possible to let $\gamma_{2}>1$. However, such a choice can reduce the convergence speed and importantly make the algorithm less robust to DoS attacks [27]. Therefore, $\gamma_{2}=1$ is the optimal value corresponding to the player dynamics in this paper. If the players’ dynamics follows $x_{i}(k+1)=Ax_{i}(k)+Bu_{i}(k)$ with $A$ having at least one unstable eigenvalues, one should let the zooming-out factor $\gamma_{2}>1$ to prevent quantizer saturation.
@@
Convergence speed to the NE majorly depends on the choices of zooming-in parameter $\gamma_{1}$ and the amount of DoS attacks. First, fix $\delta$ such that $\rho(\bar{H})$ is minimized. Subsequently, choose a small $\gamma_{1}$ satisfying $\rho(\bar{H})<\gamma_{1}<1$. By (IV-B), one can find that a small $\gamma_{1}$ can increase the convergence speed. Note that a small $\gamma_{1}$ also implies a large data rate, i.e., a fast convergence speed requires more bandwidth. This is because a $\gamma_{1}$ close to $\rho(\bar{H})$ can lead to a large $\gamma$ and hence a large $C$ (see (-B)). For the influence of DoS attacks, first note that when $1/T+\Delta/\tau_{D}$ approaches 0, $T_{S}(\cdot)$ in Lemma 1 approaches $k$. Then $\gamma_{1}^{T_{S}(\cdot)}$ converges to 0 quickly. This means that a mild DoS level admits a fast convergence speed and vice versa. Overall, a large network bandwidth and mild DoS attacks essentially lead to a large average data rate, which lead to a fast convergence speed. The discussion in this remark is also verified by the subsequent simulation results.
@@
We compare our results with those in relevant papers. Compared with the works of quantized consensus, our paper involves the nonlinear function $f_{i}(\cdot)$. Therefore, the method therein for selecting the zooming-in factor $\gamma_{1}$ is not applicable and new techniques in Lemma 3 are involved. Compared with [12], we consider the presence of DoS attacks. The techniques for analysis are also different, i.e., the analysis in [12] is based on a Lyapunov approach and ours is aligned with quantized consensus in [29, 28, 26, 27]. Compared with [6], our paper additionally considers the presence of DoS attacks and limited communication bandwidth. Both our paper and [11] consider quantized distributed NE seeking problems. As significant improvements, the control architecture in our paper can achieve asymptotic convergence to the NE instead of a practical convergence in [11]. Moreover, the ranges of quantizers preventing saturation are explicitly provided in our paper, which are not provided in [11].

### v3

@@ I Introduction
Game theory has attracted tremendous attention owing to its advances in decision making. It has great potentials in the applications such as wireless communication networks, defending cyber attacks and cloud computing [1, 2]. Nash equilibrium (NE) seeking is one of the most important problems in non-cooperative games. Recently, distributed NE seeking with partial-decision information becomes a new thriving direction, in which players do not need to have direct access to the information of all the other players [3, 4].
@@
Under an unreliable network with cyber attacks, players’ actions may not reach a NE [17]. Our paper considers one of the most common cyber attacks: Denial-of-Service (DoS) attacks. The intention of DoS attackers is to maliciously jam the networks. Communication failures induced by DoS can exhibit a temporal profile quite different from the one induced by genuine packet losses. Particularly, packet dropouts induced by DoS need not follow a given class of probability distributions [18]. Therefore, the techniques relying on probabilistic arguments for random packet losses may not be applicable in the context of DoS attacks. In the past decade, control over DoS-corrupted networks has been widely studied [19, 20]. The paper [20] proposed a new distributed resilient control method to address the output regulation problem under DoS attacks, without pre-knowing the dynamics of the leader. Since DoS attacks can induce communication failures, the existing distributed NE seeking strategies and results, e.g., in [10, 5, 11, 21], may not always hold.
@@
When quantization and DoS attacks coexist, NE seeking problems are challenging because one needs to ensure that quantizers should not saturate under DoS attacks. For example, under DoS attacks, states could diverge. If a quantization scheme has only zooming-in capabilities as in [11], quantizer saturation will occur under DoS attacks. To mitigate quantizer saturation problems in the presence of random or malicious packet losses, one can develop quantization strategies with both zooming-in and zooming-out capabilities, for centralized systems [22] and multi-agent systems [23]. For instance, the work [22] developed zooming-in and out quantization strategies to stabilize centralized systems under random packet losses. In [23], the authors proposed a zooming-in and out strategy to tackle quantized consensus problems under DoS attacks. The approach in our paper is inspired by the ideas of consensus based NE seeking in [5], dynamic quantized consensus under DoS attacks in [23] and quantized consensus in the seminal papers [24, 25].
@@
Our paper deals with quantized distributed NE seeking problems under DoS attacks. We are interested in three questions: 1. how to design a new quantization scheme realizing NE seeking and preventing quantizer saturation under DoS attacks; 2. how robust is the quantized distributed NE seeking algorithm; 3. under DoS attacks, what is the lower bound of the number of quantization levels preventing quantizer saturation. Different from the Lyapunov-based approaches [10, 11], we develop a new path to answer the above questions, which is aligned with the approaches for studying quantized consensus [24, 25, 23]. We mention that such a new development is challenging from the technical viewpoint because NE seeking problems involve nonlinear functions, under which some key techniques for linear-system consensus in [24, 25, 23] are not applicable. The main contributions of this paper are listed as follows:
@@ II-B Time-constrained DoS
For more discussions regarding Assumptions 3 and 4, we refer the readers to [26] and [29]. Note that Assumptions 3 and 4 were proposed in [29] investigating centralized systems. Later, they have been widely adopted in the works of multi-agent systems, e.g., [30, 31].
@@
[23] Consider the DoS attacks characterized by Assumptions 3 and 4 and the network sampling period $\Delta$. If $11/T+\Delta/\tau_{D}<$ 1, then $T_{S}(0,k\Delta))\geq\left(1-1/T-\Delta/\tau_{D}\right)k-\frac{\kappa+\eta\Delta}{\Delta}.$  $\blacksquare$
@@ II-C Quantizer
The limitation of bandwidth implies that transmitted signals are subject to quantization. Let $b\in\mathbb{R}$ be the original scalar signal before quantization and $q_{R}(\cdot)$ be the quantization function for a scalar input value as
@@ II-D Control objectives
    Design a quantized distributed NE seeking controller and find a $R$ (or equivalently $(2R+1)/2$) such that the quantizer (8) is not saturated at all $k$, i.e., $|b|\leq(2R+1)/2$.
@@ III-A Control architecture
with $\theta(0)=\theta_{0}\in\mathbb{R}_{>0}$. In this paper, we select the zooming-in factor $\gamma_{1}<1$ and the “zooming-out factor” $\gamma_{2}=1$. Since $\gamma_{2}=1$, “zooming-out” also implies “holding”. The reasons of such $\gamma_{1}<1$ for zooming-in and $\gamma_{2}=1$ for holding will be specified later in Remark 3. Please note that acknowledgments are not necessary because each player is able to passively know the status of DoS at $k$ by whether receiving the neighbors’ transmissions. Thus, $\theta(k)$ can be updated locally. Since each player has the identical initial condition $\theta_{0}$, they can also locally check $k\in H_{q}$ or $k\notin H_{q}$, one can verify that $\theta(k)$ is synchronized in all the players.
In (17), the update of $\theta(k)$ has two modes. The first mode is for NE seeking and the second is particularly for mitigating the influence of DoS attacks. Under DoS attacks, $y_{ij}(k+1)-\hat{y}_{ij}(k)$ and $x_{j}(k+1)-\hat{x}_{j}(k)$ in (13) and (14) can diverge [23]. If $\theta(k)$ always zooms in by only utilizing $\gamma_{1}$ as in [11], then as $\theta(k)$ approaches zero, $|(y_{ij}(k+1)-\hat{y}_{ij}(k))/\theta(k)|$ and $|(x_{j}(k+1)-\hat{x}_{j}(k))/\theta(k)|$ in $q_{R}(\cdot)$ can approach infinity. Eventually, quantizer saturation will occur, e.g., $|(y_{ij}(k+1)-\hat{y}_{ij}(k))/\theta(k)|>(2R+1)/2$. Quantizer saturation will also occur in general if the players transmit quantized $y_{ij}(k+1)/\theta(k)$ to other players instead of following the mechanism in (13), even without DoS. This is because the growing $y_{ij}(k+1)/\theta(k)$ (due to $\theta(k)\to 0$) will exceed the range of any finite quantizer $q_{R}(\cdot)$. The same problem also exists for (14).
@@
By (11)–(17), one can see that the NE seeking controller has two modes depending on the status of DoS attacks. We propose such a switched-type controller for mitigating the influence of DoS attacks. The laws for quantized NE seeking in (12)–(17) are partially inspired by the approaches in quantized consensus [25, 24, 23]. However, as will be shown later, due to the presence of nonlinear function in (11), the detailed technical analysis is different from quantized consensus.
@@ III-B Quantized system dynamics
Case 1) To obtain (22), it is important to note the following dynamics for Case 1): $u_{i}(k)=-\delta\frac{\partial f_{i}}{\partial x_{i}}(x_{i}(k),y_{-i}(k))$ by (11) and hence $x_{i}(k+1)=x_{i}(k)-\delta\frac{\partial f_{i}}{\partial x_{i}}(x_{i}(k),y_{-i}(k))$, $\sigma(k)=h$ by (12b) and hence $y_{ij}(k+1)=y_{ij}(k)-\sigma(k)(\sum_{l=1}^{N}a_{il}(\hat{y}_{ij}(k)-\hat{y}_{lj}(k))+a_{ij}(\hat{y}_{ij}(k)-\hat{x}_{j}(k)))$ by (12a), $\hat{y}_{ij}(k+1)=\hat{y}_{ij}(k)+\theta(k)\hat{Q}_{ij}(k+1)$ by (13a), $\hat{x}_{j}(k+1)=\hat{x}_{j}(k)+\theta(k)\hat{Q}_{j}(k+1)$ by (14a). Then, one can obtain the iterative equations of $\bar{y}(k)$, $x(k)$, $e^{x}(k)$ and $e^{y}(k)$ by the dynamics above and the definitions in (18)–(21):
@@
Case 2) One should note that: $u_{i}(k)$, $x_{i}(k+1)$, $y_{ij}(k+1)$ follow the same dynamics as in Case 1) because of $k\notin H_{q}$. However, one has $\hat{y}_{ij}(k+1)=\hat{y}_{ij}(k)$ by (13a) and $\hat{x}_{j}(k+1)=\hat{x}_{j}(k)$ by (14a) because of $k+1\in H_{q}$. Then, one has
@@
$e^{y}(k+1)=Ge^{y}(k)-hA_{0}(\mathbf{1}_{N}\otimes e^{x}(k))-hS\bar{y}(k).$ (24d)
Case 3) To obtain (25), one has $u_{i}(k)=0$ by (11) and hence $x_{i}(k+1)=x_{i}(k)$, $\sigma(k)=0$ and hence $y_{ij}(k+1)=y_{ij}(k)$ by (12) because of $k\in H_{q}$; $\hat{y}_{ij}(k+1)$ and $\hat{x}_{j}(k+1)$ follow the same dynamics as in Case 2) because of $k+1\in H_{q}$. Then, we obtain
@@
$e^{y}(k+1)=e^{y}(k).$ (25d)
Case 4) To obtain (26), first note that $x_{i}(k+1)=x_{i}(k)$ and $y_{ij}(k+1)=y_{ij}(k)$ due to $k\in H_{q}$ (see Case 3). Moreover, one should also note that $\hat{y}_{ij}(k+1)=\hat{y}_{ij}(k)+\theta(k)\hat{Q}_{ij}(k+1)$ by (13a), $\hat{x}_{j}(k+1)=\hat{x}_{j}(k)+\theta(k)\hat{Q}_{j}(k+1)$ by (14a) due to $k+1\notin H_{q}$. Then, one has
@@
$e^{y}(k+1)=e^{y}(k)-\theta(k)Q_{R}\left(\frac{e^{y}(k)}{\theta(k)}\right).$ (26d)
It is clear that if the network is free from DoS attacks, Case 1) is sufficient for analyzing the quantized distributed NE seeking problem. The presence of DoS attacks complicates the NE seeking problem by introducing Cases 2)-4). Consequently, the analysis becomes more challenging due to Cases 2)-4) as shown in the proofs of Lemma 3 and Theorem 1 later. Compared with quantized consensus of linear systems, the technical challenge is raised by the involvement of the nonlinear function $P(\eta(k))$.
@@ IV-A Analysis of zooming-in and holding quantization
We follow the approach aligned with that for quantized consensus under data rate constraints [25, 24, 23] and define
@@
Note that a similar assumption on the initial condition is also required in the literature of quantized consensus [24, 25]. For quantized NE seeking problems, one also needs the assumption $C_{x^{*}}\geq|x^{*}_{i}|$. They both are required for calculating an upper bound of players’ actions and preventing quantizer saturation at the initial time. We refer the readers to Remark 1 in [11] for more detailed reasons.
@@
where $l:=\max\{l_{i}\}_{i\in\mathcal{N}}$ (given in Assumption 1). If $\|\xi^{y}(s_{g})\|_{\infty}\leq 1/2\gamma_{1}$ and $\|\xi^{x}(s_{g})\|_{\infty}\leq 1/2\gamma_{1}$ for $g=-1,0,\cdots,r-1$, then
@@
2) The purposes of Lemma 3 are two-fold: Specify the zooming-in parameter $\gamma_{1}$ and provide an upper-bound on $\|\pi(s_{r})\|$. Note that $10<\gamma_{1}<$ 1 is necessary for the success of NE seeking even without DoS attacks. The result $\|\pi(s_{r})\|<C$ is also necessary since we will use it to show that $\|\pi(\cdot)\|$ is upper bounded for all $k$. Then one can obtain the main results: Quantizer unsaturation at all $k$ and NE seeking as $k\to\infty$.
@@ IV-B Quantizer unsaturation and convergence to the NE
Proof. First, we present the basic ideas of the proof. For result i), its proof is conducted by induction: If the quantizer is not saturated at $s_{p}$ ($p=-1,0,\cdots,r$, where DoS is present at $k=0$; $p=0,\cdots,r$, where DoS is absent at $k=0$), then it will not saturate for the transmissions at $k\in(s_{r},s_{r+1}]$. This also manifests that the quantizer is not saturated at $s_{r+1}$, which implies that the quantizer will not saturate for $k\in(s_{r+1},s_{r+2}]$. By induction, one can verify that the quantizer is not saturated at any $k$ in $s_{l}\bigcup_{r\in\mathbb{Z}_{\geq l}}(s_{r},s_{r+1}]$ ($l=-1$ if DoS is present at $k=0$; $l=0$ if DoS is absent at $k=0$), which are essentially all the transmission attempts. For obtaining result ii), it is sufficient to show that $\|\pi(k)\|$ is upper bounded and the scaling function $\theta(k)\to 0$ by the fact $\left\|\!\begin{bmatrix}\|\bar{y}(k)\|\\ \|x(k)-x^{*}\|\end{bmatrix}\!\right\|=\theta(k)\|\pi(k)\|$. In particular, the scaling function $\theta(k)\to 0$ if DoS attacks satisfy $11/T+\Delta/\tau_{D}<$ 1.
@@
where the first inequalities in (IV-B) and (IV-B) are obtained by Case 4 in Section III.
@@
Step 5. In the following, we prove result ii). By (27) and (-B) (in Appendix-A), one can verify that $\|\pi(s_{r}+1)\|$ is upper bounded. Then by the first two equations in Cases 3) and 4), one can obtain $\|\pi(s_{r}+m)\|=\|\pi(s_{r}+1)\|$ for $m=2,3,\cdots$. Thus, one can see that $\|\pi(\cdot)\|$ is upper bounded at $s_{r}$ and also $s_{r}+m<s_{r+1}$ ($m=1,2,\cdots$), which implies that $\|\pi(\cdot)\|$ is upper bounded at all $k$. By the transformation in (27) and the scaling function in (17), one can obtain
@@
By Theorem 1, it is clear that if the amount of DoS attacks characterized by $1/T+\Delta/\tau_{D}$ is smaller than 1, the actions of players converge to the NE under quantized signals. As indicated in [26], this is “the maximum resilience” one can achieve in the sense that if $1/T+\Delta/\tau_{D}\geq 1$, DoS attacks can be constantly present or fully synchronized with all the transmission attempts. If the communication links among the players are totally disabled, it is impossible to solve the distributed NE seeking problem for any controller, to the best of our knowledge.
@@
Our result regarding zooming-out in dynamic quantization is consistent with those of quantized consensus [23] and quantized stabilization of centralized systems [22]. In [23], the agents have a general form (e.g., $x_{i}(k+1)=Ax_{i}(k)+Bu_{i}(k)$) and hence the zooming-out factor $\gamma_{2}$ in [23] depends on the unstable eigenvalues of $A$. In our paper, the players take the form of an integrator or equivalently $A$, as a special case, is an identity matrix. Then the zooming out factor $\gamma_{2}$ takes the value 1. It is also possible to let $\gamma_{2}>1$. However, such a choice can reduce the convergence speed and importantly make the algorithm less robust against DoS attacks [23]. Therefore, $\gamma_{2}=1$ is the optimal value corresponding to the player dynamics in this paper. If the players’ dynamics follows $x_{i}(k+1)=Ax_{i}(k)+Bu_{i}(k)$ with $A$ having at least one unstable eigenvalues, one should let the zooming-out factor $\gamma_{2}>1$ to prevent quantizer saturation.
@@
Convergence speed to the NE majorly depends on the choice of zooming-in parameter $\gamma_{1}$ and the amount of DoS attacks. First, fix $\delta$ such that $\rho(\bar{H})$ is minimized. Subsequently, choose a small $\gamma_{1}$ satisfying $\rho(\bar{H})<\gamma_{1}<1$. By (IV-B), one can find that a small $\gamma_{1}$ can increase the convergence speed. Note that a small $\gamma_{1}$ also implies a large data rate, i.e., a fast convergence speed requires more bandwidth. This is because a $\gamma_{1}$ close to $\rho(\bar{H})$ can lead to a large $\gamma$ and hence a large $C$ (see (-B) and (-B)). For the influence of DoS attacks, first note that when $1/T+\Delta/\tau_{D}$ approaches 0, $T_{S}(\cdot)$ in Lemma 1 approaches $k$. Then $\gamma_{1}^{T_{S}(\cdot)}$ converges to 0 quickly. This means that a mild DoS level admits a fast convergence speed and vice versa. Overall, a large network bandwidth and mild DoS attacks essentially lead to a large average data rate, which lead to a fast convergence speed. The discussion in this remark is also verified by the subsequent simulation results.
@@
We compare our results with those in relevant papers. Compared with the works of quantized consensus, our paper involves the nonlinear function $f_{i}(\cdot)$. Therefore, the method therein for selecting the zooming-in factor $\gamma_{1}$ is not applicable and new techniques in Lemma 3 are involved. Compared with [11], we consider the presence of DoS attacks. The techniques for analysis are also different, i.e., the analysis in [11] is based on a Lyapunov approach and ours is aligned with quantized consensus in [25, 24, 23]. Compared with [5], our paper additionally considers the presence of DoS attacks and limited communication bandwidth. Both our paper and [10] consider quantized distributed NE seeking problems. As significant improvements, the control architecture in our paper can achieve asymptotic convergence to the NE instead of a practical convergence in [10]. Moreover, the ranges of quantizers preventing saturation are explicitly provided in our paper, which are not provided in [10].